Data on college graduates' employment indicate that 35% of the graduates find a part-time job within a year of their graduation, 25% of them find a full-time job and the rest of them are unemployed.
Round your final answer to 1 decimal place and express your answer in percent form (i.e. 30.0% instead of 0.3)
If we randomly select 2 graduate students within a year of their graduation, what is the probability that at least one of them has a job (full-time or part-time)?

Answer by ikleyn(52879)

.

The probability that one single college graduate student has a job  within a year of the graduation is 0.35+0.25 = 0.6.


The probability that of 2 randomly selected college graduate students at least one of them has a job is

     P = 0.6 + 0.6 - 0.6*0.6 = 0.84.    ANSWER



Use the formula  P(at least one) = P(one) + P(other) - P(both).


I assume also that getting the job for two students are independent events.

Round as you want (or as you need).